# Local Lesion Generation is Effective for Capsule Endoscopy Image Data Augmentation in a Limited Data Setting

### v1

@@ Abstract
Limited medical imaging datasets challenge deep learning models by increasing risks of overfitting and reduced generalization, particularly in Generative Adversarial Networks (GANs), where discriminators may overfit, leading to training divergence. This constraint also impairs classification models trained on small datasets. Generative Data Augmentation (GDA) addresses this by expanding training datasets with synthetic data, although it requires training a generative model. We propose and evaluate two local lesion generation approaches to address the challenge of augmenting small medical image datasets. The first approach employs the Poisson Image Editing algorithm, a classical image processing technique, to create realistic image composites that outperform current state-of-the-art methods. The second approach introduces a novel generative method, leveraging a fine-tuned Image Inpainting GAN to synthesize realistic lesions within specified regions of real training images. A comprehensive comparison of the two proposed methods demonstrates that effective local lesion generation in a data-constrained setting allows for reaching new state-of-the-art results in capsule endoscopy lesion classification. Combination of our techniques achieves a macro F1-score of 33.07%, surpassing the previous best result by 7.84 percentage points (p.p.) on the highly imbalanced Kvasir Capsule Dataset, a benchmark for capsule endoscopy. To the best of our knowledge, this work is the first to apply a fine-tuned Image Inpainting GAN for GDA in medical imaging, demonstrating that an image-conditional GAN can be adapted effectively to limited datasets to generate high-quality examples, facilitating effective data augmentation. Additionally, we show that combining this GAN-based approach with classical image processing techniques further enhances the results.
@@ 1 Introduction
Over the past four decades, numerous methods have been developed to address issues related to low data availability and class imbalance. These include sampling techniques [34, 46, 85, 18, 26], transfer learning [68, 82], and data augmentation strategies [93, 51], particularly Generative Data Augmentation (GDA) [109, 101, 19, 8, 20]. Training effective models in a limited data regime is complicated.
The most common use case for GDA is de novo generation using deep generative modes like GANs, VAEs, and diffusion models [12, 15, 86]. These models learn the data distribution from a training set and generate synthetic labeled examples to augment the original dataset. Empirical [101, 19, 8, 20] and theoretical [109] evidence shows that GDA improves model generalization, especially when overfitting occurs.
@@
We propose IIDA, a novel data augmentation approach using a fine-tuned image inpainting model, particularly effective in low-data settings. This new approach leverages the abundance of healthy tissue, common in many medical imaging domains, to generate lesions.
@@
We compare Poisson Blending, a classical image processing method, with deep generative models for augmenting local pathological changes, showing that both techniques can enhance machine learning models. IIDA shows better performance than PBDA as a standalone data augmentation technique, however their combination provides the best results.
@@ 2.2 Image generation in capsule endoscopy
Diamantis et al. [29] introduced a VAE model trained on a subset of Kvasir Capsule and another VCE dataset. They trained distinct models for normal and abnormal image generation, achieving high-quality synthetic images. Xiao et al. [99] proposed a GAN architecture to generate synthetic capsule endoscopy images from Kvasir Capsule data, which were then used in GDA for object detection. Unlike Xiao et al., who utilized a random data split, our approach adopts the official data split, enhancing comparability with other studies and better simulating clinical environments.
@@ 2.3 Image inpainting
Early approaches to inpainting relied on patch-matching techniques [25, 10] and partial-differential equation (PDE)-based methods [17]. The introduction of GANs marked a significant shift, as researchers moved towards these models for more sophisticated inpainting tasks.
@@
Liu et al. [54] performed one of the first high-quality image inpainting for irregular holes using deep learning. Authors proposed Partial Convolution which conditions the network on valid pixels to inpaint masked region.
Subsequent advancements addressed limitations in Partial Convolution. Yu et al. [104] developed Gated Convolution, a method that learns to condition both masked and unmasked regions within an image, enhancing the model’s adaptability. They also introduced SN-PatchGAN, a GAN architecture designed to discriminate between image patches. This model allowed for user guidance in inpainting, utilizing sketches, lines, dots, and other structures as input to guide the inpainting process.
Suvorov et al. [81] utilized Fast Fourier Convolutions (FFC) [22] to develop a class of Large Mask Inpainting (LaMa) models. They showed that a large receptive field is essential for high-quality inpainting of large masks. FFC significantly enhances LaMa’s ability to recreate repetitive patterns. Additionally, Suvorov et al. proposed the High Resolution Field Perceptual Loss, leveraging a segmentation network to extract high-level features and enhance the model’s focus on object structure over texture.
Zhao et al. [108] studied co-modulation and conditional modulation in detail. Their findings demonstrated that co-modulation with a mapping network improves stochastic generative capabilities. Conditioning the synthesis network on an encoder’s latent representation and a mapped random vector enabled their model to produce realistic content, even when working with large missing areas and limited conditional information.
@@ 2.4 Generative data augmentation
Tao [83] proposed a Denoising Diffusion Probabilistic Model (DDPM)-based GDA method for surface defect inspection in a limited data regime. Synthetic samples are generated by conditioning the model on partially erased real defect samples.
@@ 2.5 Image composition
It has been used in data augmentation to improve downstream task performance in medical imaging. Wang et al. [93] utilized Poisson Blending to augment diabetic retinopathy lesions in fundus images from retinography. Lee and Cho [51] created an augmentation pipeline using Poisson Blending to augment training sets in classical endoscopy. Their solution involved the use of Grad-CAM [78] for automatic segmentation of lesions in images. They use the lowest RGB variance sliding window search to find a suitable area for blending.
@@ 3.1 Poisson Blending Data Augmentation
In this work, to enhance data augmentation effectiveness, we designed the Poisson Blending Data Augmentation pipeline, consisting of the following steps:
Image deduplication with feature extraction.
Identifying the blending location.
Performing image blending.
The method is illustrated in Figure 2 and each of these steps will be explained in detail in the following sections. While Poisson Blending has been applied in recent studies as a data augmentation technique [51, 93], our proposed pipeline is novel and specifically tailored for Capsule Endoscopy data, although it can be adapted to other medical imaging fields. The next section will elaborate on the processes of deduplication and identifying the blending location.
3.1.1 Image deduplication
A common challenge with Capsule Endoscopy data is the presence of numerous seemingly duplicate images. This occurs when the capsule remains stationary in the intestine for extended periods, often due to slow or infrequent peristaltic movements. Removing these redundant images increases dataset diversity, which enhances the variety of pathological and healthy tissue pairings in the proposed pipeline, ultimately leading to a more diverse synthetic dataset.
The deduplication process, detailed in Algorithm 1, begins by extracting latent vectors from all images in the training dataset using the DinoV2-Giant model [65]. The dataset is then filtered iteratively, removing images that are most similar to the one currently sampled. Similarity is measured by the Euclidean distance between latent vectors, with a threshold determined experimentally.
3.1.2 Identifying the blending location
To explain our approach to selecting an optimal blending location, we first introduce new definitions. Similar to Section 3.1, let represent an image domain, defined as a closed subset of and let be a closed subset of , where the previously introduced function is defined over . Additionally, a correction function , defined on , is such that [70], which will help in understanding Poisson interpolation.
@@ 3.1.2 Identifying the blending location
The function represents the adjustments needed for the pathology to blend seamlessly using the Poisson Blending algorithm. Therefore for the pathology to remain consistent after the blending, the difference must be minimal along the boundary . If this difference is too large, it can cause color bleeding or blending artifacts [3], potentially altering the semantics (or class) of and introducing label noise into the synthesized dataset. Our pipeline for finding pairs of and values consists of two steps: first, finding the most similar pair of images defined over and , and then identifying the optimal location for seamless cloning.
Identifying the most similar pair of source and target images relies on semantic search within the latent space of the DinoV2-Giant model [65]. The process is detailed in Algorithm 2. To identify target pairs for a given source, simply most similar normal images are chosen for a given lesion. After selecting such pairs, choosing an appropriate blending location becomes essential. Not all areas in Capsule Endoscopy images are suitable for seamlessly blending lesions. Regions like edges, folds, or holes are generally unsuitable because the source lesion cannot appear there without significant alteration. The Poisson Blending algorithm integrates the lesion as is, adjusting only the color for seamless blending but not modifying its texture or shape. Furthermore, if there is a large gradient difference at the boundary of the lesion and the new region (e.g., near folds or dark zones), blending artifacts may arise.
To blend a pathological subregion from a source image into a target image, it is necessary to identify the region of interest (ROI) in the target image that minimizes color differences between the two images, particularly along the border pixels. This process is illustrated in Figure 2.
Formally, let denote the set of border pixels for the bounding box in the source image, and represent the border pixels for the bounding boxes of each potential ROI in the target image, indexed by , where is the total number of ROIs generated using a sliding window algorithm. For all , the number of border pixels is denoted as .
@@
where are the border pixels of the source image, and are the corresponding border pixels in the target image for each ROI .
@@ 3.2 Image Inpainting Data Augmentation
The proposed solution leverages an Image Inpainting model, LaMa [81], to generate new structures at selected locations for data augmentation. In this work, the LaMa model was fine-tuned on a subset of the Kvasir Capsule Dataset [79] to achieve state-of-the-art data augmentation on this small dataset.
To the best of our knowledge, this is the first method to use a pre-trained Image Inpainting model, fine-tuned to generate new objects for data augmentation in classification tasks on a limited medical imaging dataset. The following sections will describe the LaMa model, the preparation of training datasets, and the model’s fine-tuning process.
3.2.1 LaMa
Image inpainting aims to fill in missing regions of an image in a visually convincing way. Following notation introduced by Suvorov et al. [81], the mask is represented as , and the masked image is . The LaMa model takes as input a stack of the masked image and the mask, denoted . The inpainting network, referred to as the generator or synthesis network, is denoted by , where are the parameters of the network. The LaMa model processes using a fully convolutional approach, yielding a three-channel RGB image . Training involves preparing image-mask pairs.
The LaMa inpainting network features an architecture (Figure 3) similar to fully convolutional ResNet with 3 downsampling blocks, 6-18 residual blocks, and 3 upsampling blocks. The discriminator network is based on PatchGAN [42]. For residual blocks, Suvorov et al. [81] use Fast Fourier Convolution (FFC) [22]. FFCs are fully differentiable and can effectively replace standard convolutions in the architecture. This enables the network to receive global context information already in the first stages, enhancing its content generation capabilities in situations where limited conditional information is available, such as when large masks are applied. Additionally, the use of FFCs improves the network’s ability to identify and reconstruct repetitive patterns, which should be beneficial in the Capsule Endoscopy data domain.
Suvorov et al. [81] introduced High Receptive Field Perceptual Loss (HRFPL). It is based on perceptual loss from [43], which computes the distance between feature vectors extracted from a pretrained network applied to generated and original images. Suvorov et al. argue, that for large mask inpainting, it is important to understand the global structure from early layers also in the pretrained network for HRFPL loss. Therefore, the base model is implemented with Fourier or Dilated convolutions. The formula for HRFPL loss, as defined by [81], is provided below:
@@ 3.2.1 LaMa
LaMa uses adversarial loss based on image patches. Only patches that intersect with the masked region of interest are considered ”fake”. The authors decided to use non-saturating adversarial loss, defined in [81] as:
where is a sample from dataset, is a synthetically generated mask, is the inpainting result for , stops gradients w.r.t and is the joint loss to optimize.
@@ 3.2.2 Fine-tuning
It has been shown in the literature that GANs fine-tuning is possible [60, 63, 97, 98]. LaMa utilizes training-time data augmentation for discriminator, which has been shown by Karras et al. [45] to be effective when training on a limited dataset.
We tried to train the Big LaMa-Fourier model from scratch on the Kvasir Capsule Dataset, using the training setup described in [81]. The model could not learn to generate plausible content in the capsule endoscopy images when trained from scratch. Therefore in this work, the Big LaMa-Fourier model trained by Suvorov et al. on Places2 Dataset has been fine-tuned using a modified loss function:
where . Adding the L1 term between the reconstruction of lesion generated by the model, and the original pathology, with , makes the reconstruction task easier for the model, while other parts of the loss function, such as help to leverage the problem of blurriness when applying L1 loss term. This term is introduced to ensure that the model does not deviate from what it learned prior to the fine-tuning process. The rest of the scaling factors, are set with the values used by [81]. The LaMa model, as presented by Suvorov et al. [81], does not allow for class conditioning and the development of such methods is out of the scope of this paper. Therefore, a separate model is fine-tuned for each of the considered lesion classes. In the following paragraphs, the fine-tuning process of separate models for each class will be described.
Formally, let be a subset of Euclidean space corresponding to image dimensions, representing capsule endoscopy images, and be the set of corresponding labels, where represents the healthy tissue and represent different lesion classes. Let be a dataset of observations, where and .

### v2

@@ Abstract
Limited medical imaging datasets challenge deep learning models by increasing risks of overfitting and reduced generalization, particularly in Generative Adversarial Networks (GANs), where discriminators may overfit, leading to training divergence. This constraint also impairs classification models trained on small datasets. Generative Data Augmentation (GDA) addresses this by expanding training datasets with synthetic data, although it requires training a generative model. We propose and evaluate two local lesion generation approaches to address the challenge of augmenting small medical image datasets. The first approach employs the Poisson Image Editing algorithm, a classical image processing technique, to create realistic image composites that outperform current state-of-the-art methods. The second approach introduces a novel generative method, leveraging a fine-tuned Image Inpainting GAN to synthesize realistic lesions within specified regions of real training images. A comprehensive comparison of the two proposed methods demonstrates that effective local lesion generation in a data-constrained setting allows for reaching new state-of-the-art results in capsule endoscopy lesion classification. Combination of our techniques achieves a macro F1-score of 33.07%, surpassing the previous best result by 7.84 percentage points (p.p.) on the highly imbalanced Kvasir Capsule Dataset, a benchmark for capsule endoscopy. To the best of our knowledge, this work is the first to apply a fine-tuned Image Inpainting GAN for GDA in medical imaging, demonstrating that an image-conditional GAN can be adapted effectively to limited datasets to generate high-quality examples, facilitating effective data augmentation. Additionally, we show that combining this GAN-based approach with classical image processing techniques further improves the results.
@@ 1 Introduction
Over the past four decades, numerous methods have been developed to address issues related to low data availability and class imbalance. These include sampling techniques [34, 46, 85, 18, 26], transfer learning [68, 82], and data augmentation strategies [93, 51], particularly Generative Data Augmentation (GDA) [109, 101, 19, 8, 20]. However, training effective models in a limited data regime remains complicated.
The most common use case for GDA is de novo generation using deep generative models like GANs, VAEs, and diffusion models [12, 15, 86]. These models learn a data distribution from a training set and generate synthetic labeled examples to augment the original dataset. Empirical [101, 19, 8, 20] and theoretical [109] evidence shows that GDA improves model generalization, especially when overfitting occurs.
@@
We propose IIDA method, a novel data augmentation approach using a fine-tuned image inpainting model, particularly effective in low-data settings. This new approach leverages the abundance of healthy tissue, common in many medical imaging domains, to generate lesions.
We designed PBDA method, a data augmentation pipeline adjusted to specific characteristics of VCE data. This pipeline uses image processing algorithms to generate high-quality synthetic lesions by modifying images of healthy tissue. Unlike IIDA, PBDA can be used with any dataset size.
@@
We compare PBDA with IIDA for augmenting local pathological changes, showing that both techniques can enhance machine learning models. IIDA shows better performance than PBDA as a standalone data augmentation technique, however their combination provides the best results.
@@ 2.2 Image generation in capsule endoscopy
Diamantis et al. [29] introduced a VAE model trained on a subset of Kvasir Capsule and another VCE dataset. They trained distinct models for normal and abnormal image generation, achieving high-quality synthetic images. Xiao et al. [99] proposed a GAN architecture to generate synthetic capsule endoscopy images from Kvasir Capsule data, which were then used in GDA for object detection. Unlike Xiao et al. [99], who utilized a random data split, our approach adopts the official data split, enhancing comparability with other studies and better simulating clinical environments.
@@ 2.3 Image inpainting
Suvorov et al. [81] utilized Fast Fourier Convolutions (FFC) [22] to develop a class of Large Mask Inpainting (LaMa) models. They showed that a large receptive field is essential for high-quality inpainting of large masks. FFC significantly enhances LaMa’s ability to recreate repetitive patterns. Additionally, Suvorov et al. [81] proposed the High Resolution Field Perceptual Loss, leveraging a segmentation network to extract high-level features and enhance the model’s focus on object structure over texture.
@@ 2.4 Generative data augmentation
Tao et al. [83] proposed a Denoising Diffusion Probabilistic Model (DDPM)-based GDA method for surface defect inspection in a limited data regime. Synthetic samples are generated by conditioning the model on partially erased real defect samples.
@@ 2.5 Image composition
It has been used in data augmentation to improve downstream task performance in medical imaging. Wang et al.[93] utilized Poisson Blending to augment diabetic retinopathy lesions in fundus images from retinography. Lee et al. [51] created an augmentation pipeline using Poisson Blending to augment training sets in classical endoscopy. Their solution involved the use of Grad-CAM [78] for automatic segmentation of lesions in images. They use the lowest RGB variance sliding window search to find a suitable area for blending.
@@ 3 Methods
This section introduces the PBDA and IIDA methods developed in this study, with a clear presentation of each method’s pipeline, background, and data preparation steps. A visual summary of the methods is provided in Figure 1.
In Section 3.1, we outline the PBDA pipeline, beginning with background information on the Poisson Blending (Section 3.1.1) and detailing the data preparation steps (Section 3.1.2). These data preparation techniques improve data augmentation outcomes and are used consistently with both PBDA and IIDA to enable a direct comparison between the two approaches.
Section 3.2 then presents the IIDA, beginning with background information on LaMa (Section 3.2.1), followed by the finetuning process (Section 3.2.2), and concluding with the data generation process (Section 3.2.3). This structured approach ensures a coherent understanding of PBDA and IIDA, emphasizing the methods’ individual contributions and shared comparative framework.
@@ 3.1 Poisson Blending Data Augmentation
To optimize data augmentation for VCE images, we developed a Poisson Blending Data Augmentation pipeline that increases data diversity while addressing high redundancy of VCE data. The method aims to create augmented samples that appear natural, while introducing variability. Our pipeline incorporates the following stages:
Feature extraction.
Image deduplication.
Image pairs selection.
Region of interest (ROI) selection.
Poisson Blending.
The method is illustrated in Figure 1 and each of these steps will be explained in detail in the following sections. While Poisson Blending has been applied in recent studies as a data augmentation technique [51, 93], our proposed pipeline is novel and specifically tailored for VCE data, although it can be adapted to other medical imaging fields.
To support understanding of our methodology, the following section provides background on Poisson Blending, laying the groundwork for the data preparation procedure described in Section 3.1.2.
3.1.1 Background on Poisson Blending
@@
3.1.2 Data preparation
Our data preparation procedure (Figure 3) utilizes specific characteristics of VCE data and the Poisson Blending technique to maximize the data augmentation effectiveness. It consists of feature extraction, image deduplication, and image pairs and ROI selection steps, which are further described in detail.
Feature extraction and image deduplication. A common challenge with VCE data is the presence of numerous seemingly duplicate images. This occurs when the capsule remains stationary in the intestine for extended periods, often due to slow or infrequent peristaltic movements. Removing these redundant images increases dataset diversity, which enhances the variety of pathological and healthy tissue pairings in the proposed pipeline, ultimately leading to a more diverse synthetic dataset.
The deduplication process, detailed in Appendix A, begins by extracting latent vectors from all images in the training dataset using the DinoV2-Giant model [65]. The dataset is then filtered iteratively, removing images that are most similar to the one currently sampled. Similarity is measured by the Euclidean distance between latent vectors, with a threshold determined experimentally.
Image pairs selection. To augment data through Poisson Blending, firstly appropriate ROIs need to be found for lesion blending. To explain our approach to selecting an optimal blending location, we first introduce new definitions. Similar to Section 3.1.1, let represent an image domain, defined as a closed subset of and let be a closed subset of , where the previously introduced function is defined over . Additionally, a correction function , defined on , is such that [70], which will help in understanding Poisson interpolation.
@@ 3.1.2 Data preparation
The function represents the adjustments needed for the lesion to blend seamlessly using the Poisson Blending algorithm. Therefore for the lesion to remain consistent after the blending, the difference must be minimal along the boundary . If this difference is too large, it can cause color bleeding or blending artifacts [3], potentially altering the semantics (or class) of and introducing label noise into the synthesized dataset. Our pipeline for finding pairs of and values consists of two steps: first, finding the most similar pair of images defined over and , and then identifying the optimal location for seamless cloning.
Identifying the most similar pair of source and target images relies on semantic search within the latent space of the DinoV2-Giant model [65]. The process is detailed in Appendix A. To identify target pairs for a given source, simply most similar images of healthy tissue are chosen for a given lesion. After selecting such pairs, choosing an appropriate blending location becomes essential. Not all areas in VCE images are suitable for seamlessly blending lesions. Regions like edges, folds, or holes are generally unsuitable because the source lesion cannot appear there without significant alteration. The Poisson Blending algorithm integrates the lesion as is, adjusting only the color for seamless blending but not modifying its texture or shape. Furthermore, if there is a large gradient difference at the boundary of the lesion and the new region (e.g., near folds or dark zones), blending artifacts may arise.
ROI selection. To blend a pathological subregion from a source image into a target image, it is necessary to identify a ROI in the target image that minimizes the color difference to the ROI in the source image along the border pixels. This process is illustrated in Figure 3.
Formally, let denote the set of border pixels for the lesion’s bounding box in the source image, and represent the border pixels for the bounding boxes of each potential ROI in the target image, indexed by , where is the total number of ROIs generated using a sliding window algorithm. For all , the number of border pixels is denoted as .
@@
where are the border pixels of ROI in the source image, and are the corresponding border pixels in the target image for each ROI .
@@ 3.2 Image Inpainting Data Augmentation
The IIDA method leverages the unique strength of image inpainting — its capacity to fill missing image regions in a visually convincing manner — to generate realistic lesions in images of healthy tissue. By harnessing this feature, IIDA enhances synthetic sample generation and enables effective data augmentation. The approach involves training an image inpainting model to reconstruct masked-out lesions, conditioned on surrounding tissue. This conditioning directs the model to sample lesions that are better aligned with the geometry, brightness, contrast, texture, and shape of the background tissue.
IIDA requires selecting a backbone model for image inpainting; in this study, LaMa [81] was chosen, as it demonstrated the ability to generate realistic capsule endoscopy images in a zero-shot setting (Figure 4). It was fine-tuned on a subset of the Kvasir Capsule Dataset [79] to achieve state-of-the-art data augmentation.
To the best of our knowledge, this is the first method to use a pre-trained Image Inpainting model, fine-tuned to generate new objects for data augmentation in classification tasks on a limited medical imaging dataset.
The following sections will provide background information on the LaMa model, the model’s fine-tuning, and the data generation process.
3.2.1 Background on LaMa
Following notation introduced by Suvorov et al. [81], the mask is represented as , and the masked image is . The LaMa model takes as input a stack of the masked image and the mask, denoted . The inpainting network, referred to as the generator or synthesis network, is denoted by , where are the parameters of the network. The LaMa model processes using a fully convolutional approach, yielding a three-channel RGB image . In this study, bounding boxes provided for lesions in the Kvasir Capsule Dataset [81] are used to prepare image-mask pairs.
Suvorov et al. [81] introduced High Receptive Field Perceptual Loss (HRFPL). It is based on perceptual loss from Johnson et al. [43], which computes the distance between feature vectors extracted from a pretrained network applied to generated and original images. The formula for HRFPL loss, as defined by Suvorov et al. [81], is provided below:
@@ 3.2.1 Background on LaMa
LaMa uses adversarial loss based on image patches. Only patches that intersect with the masked region of interest are considered fake. The non-saturating adversarial loss is defined by Suvorov et al. [81] as:
where stops gradients w.r.t and is the joint loss to optimize.
@@ 3.2.2 Fine-tuning
Multiple studies have shown that GANs fine-tuning is possible [60, 63, 97, 98]. LaMa utilizes training-time data augmentation for discriminator, which has been shown by Karras et al. [45] to be effective when training on a limited dataset. The primary novelty of this paper is a modification of the standard training procedure of image inpainting models to fine-tuning on the task of lesion reconstruction.
Therefore in this work, the Big LaMa-Fourier model trained by Suvorov et al. [81] on Places2 Dataset has been fine-tuned using a modified loss function:
where . Adding the L1 term between the reconstruction of lesion generated by the model, and the original pathology, with , makes the reconstruction task easier for the model, while other parts of the loss function, such as help to leverage the problem of blurriness when applying L1 loss term. This term is introduced to ensure that the model does not deviate from what it learned prior to the fine-tuning process. The rest of the scaling factors, are set with the values used by Suvorov et al. [81]. The LaMa model, as presented by Suvorov et al. [81], does not allow for class conditioning and the development of such methods is out of the scope of this paper. Therefore, a separate model is fine-tuned for each of the considered lesion classes. In the following paragraphs, the fine-tuning process of separate models for each class will be described.
Formally, let be a subset of Euclidean space corresponding to image dimensions, representing capsule endoscopy images, and be the set of corresponding labels, where represents the healthy tissue and represent different lesion classes. Let be a dataset of observations, where , and , such that an observation is coupled with a mask derived from a bounding box for a lesion or obtained through ROI selection procedure (Section 3.1.2) for healthy tissue.
@@
3.2.3 Data generation process
Following the formulations introduced in the previous section, a general approach to generating synthetic data using image inpainting is described in the following paragraph.